OpenStudy (anonymous):

does the summation of infinity of n=3 of 4/sqrt(4) converge or diverge? whats the value of L. Your suppose to use the ratio test but idk how

myininaya (myininaya):

\[\sum_{i=1}^{\infty} \frac{3}{4 \sqrt{4} }?\]

myininaya (myininaya):

I don't think this is what you meant